Adriana Corporation manufactures football equipment. In planning for next year, the managers want to understand the relation between activity and overhead costs. Discussions with the plant supervisor suggest that overhead seems to vary with labor-hours, machine-hours, or both. The following data were collected from last year's operations.

Month Labor-Hours Machine-Hours Overhead Costs
1 730 1,350 $ 102,700
2 715 1,402 103,743
3 690 1,521 109,869
4 745 1,454 108,371
5 775 1,591 116,123
6 755 1,581 114,410
7 735 1,389 106,927
8 730 1,306 102,159
9 710 1,449 106,487
10 790 1,546 113,075
11 675 1,298 104,426
12 710 1,618 117,866


Required:

a. Use the high-low method to estimate the fixed and variable portions of overhead costs based on machine-hours.

b. Managers expect the plant to operate at a monthly average of 1,600 machine-hours next year. What are the estimated monthly overhead costs, assuming no inflation?
Answer #1

Solution a:

Variable portion of overhead = (cost at highest level of machine hours- cost at lowest point of machine hours) / (highest machine hours - lowest machine hours)

= ($117866- $104426) / (1618-1298) = $13440/ 320 = $42 per machine hour

Fixed portion of overhead cost = total cost - variable portion = $117866 - 1618*42 = $49,910

Solution b:

Estimated monthly overhead costs = fixed cost + variable cost = $49910 + 1600*$42 = $117,110